A block of gold measure 3.54 × 7.00 × 1.25 inches. The density of gold is 19.32 g/cm3, the current price of gold is 1,291 $/troy ounce and a troy ounce is 31.10 grams. How many cm3 is this block of gold and how much is it worth?

      Volume = ______ cm3   and Block Value = ______
Answer #1

Volume = length * breadth * height

= 3.54*7.00*1.25

= 30.975 in3

1 in3 = 16.3871 cm3

30.975 in3 = 507.59 cm3

B. Mass = volume * density

= 507.59 * 19.32

= 9806.64 grams

In Troy ounce = 9806.64 / 31.10

= 315.326

Block value = 1291 * 315.326

= 407085.866 $

If you have any query please comment

If you satisfied with the solution please like it thankxx